OpenStudy (anonymous):

The sum of two numbers is 21. When the larger number is subtracted from four times the smaller number, the difference is 4. What are the two numbers?

myininaya (myininaya):

x+y=21 is from the first sentece assume x is the larger number x-4y=4 is from the second sentence

OpenStudy (anonymous):

The 2nd sentence would by 4y - x = 4 if you assume x is the larger number.

myininaya (myininaya):

yes gj blex hes right my sentence is wrong